OpenStudy (anonymous):

log subscript 6 x + log6 (x-5)=2

OpenStudy (anonymous):

Use The property : LogX+LogY=Log(x*Y) Can U do it now ?

OpenStudy (anonymous):

not sure but ill try

OpenStudy (anonymous):

Anyway ,Here is The Answer :: \[\Large Log_{6}x+Log_{6}(x-5)=2\] Using the Property that i posted : \[\Large Log_{6}(x*(x-5))=2\] Change it to Exponential: \[\Large 6^2=x^2-5x\] now factorize and solve it and get the value of x.

OpenStudy (anonymous):

lost

OpenStudy (anonymous):

where ?

OpenStudy (anonymous):

i multiplied the logs and got 0.6 put it in for x and the answers werent coming up

OpenStudy (anonymous):

Did U follow My Steps ?

OpenStudy (anonymous):

:/ yes